# Localized Graph-Based Neural Dynamics Models for Terrain Manipulation

### v2

@@ I Introduction
Terrain manipulation is essential in construction industry and outer space exploration [1, 2]. Autonomous robots that can create desired structure in the construction site by interpreting sensory data and making decision on where and how to manipulate the terrain, can greatly improve efficiency and be deployed to place where lacks human labor, e.g., other planets. Prediction is a key challenge in terrain manipulation, since robots need to predict outcomes of excavated volume and the movement of terrain shape after scooping, dumping, or pushing material. People have traditionally built approximate models using analytical expressions and geometric calculations to predict excavation outcomes. However, these models are often limited in adapting to other action types and typically require predefined parameters [3]. Simulation-based methods use terramechanics physics simulators which can handle a wider variety of situations, but these are extremely computationally expensive and still require solving difficult parameter identification problems [4, 5, 6]. Learning-based approaches have also been investigated since early work in autonomous excavation [7], but recent attention has been focused on deep learning techniques, such as image-based prediction of granular media manipulation [8]. A line of recent work has studied the use of Graph-Based Neural Dynamics (GBND), a technique that uses Graph Neural Networks (GNNs) to model system dynamics of a scene represented as set of 3D particles, with each particle representing a small volume of matter [9, 10, 11]. GBND has been applied to various nontrivial deformable object manipulation tasks [12, 13, 14, 15], and can generalize well across different viewpoints, objects and robots poses, shapes and sizes. However, all prior such works have addressed table-top scenarios and objects with bounded volume that can be observed through static cameras.
@@
In this work, we propose a framework, L-GBND, for modeling deformable terrains at fine resolutions over large spatial scales, such as a construction site in which a mobile excavator uses 3D mapping to generate a terrain representation. In particular, our method extends the GBND to handle spatially large-scale and even unbounded terrains which may be represented by graphs of millions of particles that cannot fit in GPU memory. To predict the deformation of the terrain under a candidate interaction in a computationally efficient manner suitable for planning, our method restricts the attention of dynamics prediction to a dynamically-selected local neighborhood. We introduce a learning-based 3D region-of-interest (RoI) proposer that identifies a small subset of particles from the large graph that are likely to move. Assuming particles that are outside the RoI stay static during robot-terrain interaction improves speed by orders of magnitude in addition to higher dynamics prediction accuracy. L-GBND simultaneously learns the GBND dynamics and the RoI from simulation data, and both are fine-tuned on real data. Moreover, we introduce a novel boundary feature representation to address the problem that naïvely applying GBND to the identified subgraph leads to erroneous predictions. In our experiments, we verify both the efficiency and effectiveness of L-GBND for manipulating terrains of different materials.
@@ II RELATED WORK AND DISCUSSION
Dynamics Model Learning. Physics-based models have shown effectiveness in a wide range of robotic manipulation tasks [20, 21, 22], but accurate model building requires full state estimation. To address this issue, recent advances in machine learning are applied to learn the system dynamics models directly from sensory data [23, 24, 25, 26]. The learned dynamics model generally demonstrates superior effectiveness relative to its model-free counterparts [27]. Graph-based scene representation processed by GNNs have showcased effectiveness in dynamics modeling due to the spatial relational inductive bias injected [28, 29, 9, 10, 11]. Prior work explored a wide range of applications of graph-based dynamics model, including but not limited to rigid bodies [9, 10], elastic-plastic materials [13, 12], fluids [9, 11], and granular materials [18, 11, 30]. However, the vast majority of prior work model systems that are either confined to a 3D box or atop a 2D plane, and our work demonstrates an approach to extend graph-based approaches to terrains of unbounded breadth and depth using regions of interest and RoI boundary features.
Model-Free Policy Learning. While model-free methods such as reinforcement learning (RL) and behavior cloning (BC) have achieved notable success in robotic manipulation [31, 32, 33, 34], they typically suffer from limited generalization, particularly in environments involving complex, high-dimensional, and long-horizon interactions like terrain deformation. These approaches require extensive task-specific data collection and tuning, often making them impractical for settings involving diverse or previously unseen terrain geometries. In the context of deformable terrain manipulation, the inability of model-free policies to extrapolate beyond the training distribution presents a fundamental limitation [35]. While such methods can learn reactive strategies within the training distribution, they lack an explicit mechanism for generalizing to new goals or terrain states. In contrast, dynamics model-based approaches retain predictive structure: given a learned forward model, planning can be performed zero-shot for unseen terrain states by simulating the effects of actions. This capability enables flexible and task-agnostic generalization, which is essential for terrain-aware manipulation. Because of these limitations, we do not consider model-free policies as baselines in this work. Our focus is exclusively on dynamics model learning, as it enables both general-purpose planning and scalability across diverse terrain interaction scenarios.
@@ III Localized Graph-Based Neural Dynamics
Our key innovations include: (1) learning $f$ and $g$ to minimize L-GBND memory usage and computation time while retaining high accuracy; and (2) introducing novel node features that enable learning L-GBND over arbitrary volumetric domains, ensuring that accuracy is preserved at the boundary of $R_{t}$, and that translation invariance is maintained.
@@ III-C Adaptive Fine-Grained Region of Interest Identification
Because the computation time and memory usage of GBND are closely tied to the number of particles and edges in the graph, limiting the graph’s size helps ensure efficiency. We learn an implicit function $g(p,u_{t})$ that defines the region of interest (RoI) $R_{t}(u_{t})=\{p\in\mathbb{R}^{3}\mid g(p,u_{t})\geq 0\}$.
@@ III-F Terrain Manipulation Planning with Localized GBND
We use our L-GBND in a planner for terrain manipulation tasks. Our planner selects among parameterized scooping, pushing, and dumping trajectories to minimize the difference between the predicted terrain shape and a target heightmap. For scooping and pushing, we use the penetrate-drag-scoop (PDS) trajectory (Fig. 4) used in excavation [1, 41, 7], which maintains a 0 roll angle. This yields a 9D motion parameter space. Dumping uses a 6D parameterization (dumping location, yaw angle, pre- and post-dump pitch angles). Scooping and pushing are only allowed if the scoop is empty. The planner is based on Model Predictive Path Integral (MPPI) control [42]. A batch of trajectories is sampled in each iteration. The corresponding control sequences are rolled out with the dynamics model, and the outcomes are scored based on proximity to the goal shape. The best trajectory is selected and executed, and the terrain state is updated. Pseudocode is shown in Algo. 2.
Figure 4: Sideview of a “scoop” primitive. The “push” primitive is a special case of “scoop” primitive with $\alpha=\beta=\pi/2$. The “dump” primitive is simply pivoting the scoop about the scoop-tip. Figure borrowed from [1].

### v3

@@ I Introduction
Terrain manipulation is essential in construction industry and outer space exploration [48, 15]. Autonomous robots that can create desired structure in the construction site by interpreting sensory data and making decision on where and how to manipulate the terrain, can greatly improve efficiency and be deployed to place where lacks human labor, e.g., other planets. Prediction is a key challenge in terrain manipulation, since robots need to predict outcomes of excavated volume and the movement of terrain shape after scooping, dumping, or pushing material. People have traditionally built approximate models using analytical expressions and geometric calculations to predict excavation outcomes. However, these models are often limited in adapting to other action types and typically require predefined parameters [27]. Simulation-based methods use terramechanics physics simulators which can handle a wider variety of situations, but these are extremely computationally expensive and still require solving difficult parameter identification problems [11, 4, 34]. Learning-based approaches have also been investigated since early work in autonomous excavation [33], but recent attention has been focused on deep learning techniques, such as image-based prediction of granular media manipulation [35]. A line of recent work has studied the use of Graph-Based Neural Dynamics (GBND), a technique that uses Graph Neural Networks (GNNs) to model system dynamics of a scene represented as set of 3D particles, with each particle representing a small volume of matter [16, 3, 30]. GBND has been applied to various nontrivial deformable object manipulation tasks [31, 32, 17], and can generalize well across different viewpoints, objects and robots poses, shapes and sizes. However, all prior such works have addressed table-top scenarios and objects with bounded volume that can be observed through static cameras.
@@
In this work, we propose a framework, Localized Graph-Based Neural Dynamics (L-GBND), for modeling deformable terrains at fine resolutions over large spatial scales, such as a construction site in which a mobile excavator uses 3D mapping to generate a terrain representation. In particular, our method extends the GBND to handle spatially large-scale and even unbounded terrains which may be represented by graphs of millions of particles that cannot fit in GPU memory. To predict the deformation of the terrain under a candidate interaction in a computationally efficient manner suitable for planning, our method restricts the attention of dynamics prediction to a dynamically-selected local neighborhood. We introduce a learning-based 3D region-of-interest (RoI) proposer that identifies a small subset of particles from the large graph that are likely to move. Assuming particles that are outside the RoI stay static during robot-terrain interaction improves speed by orders of magnitude in addition to higher dynamics prediction accuracy. L-GBND simultaneously learns the GBND dynamics and the RoI from simulation data, and both are fine-tuned on real data. Moreover, we introduce a novel boundary feature representation to address the problem that naïvely applying GBND to the identified subgraph leads to erroneous predictions. In our experiments, we verify both the efficiency and effectiveness of L-GBND for manipulating terrains of different materials.
Our key innovations include: (1) learning graph dynamics and RoI to minimize L-GBND memory usage and computation time while retaining high accuracy; and (2) introducing novel node features that enable learning L-GBND over arbitrary volumetric domains, ensuring that accuracy is preserved at the boundary of RoI, and that translation invariance is maintained.
@@ II RELATED WORK AND DISCUSSION
Dynamics Model Learning. Physics-based models have shown effectiveness in a wide range of robotic manipulation tasks [23, 45, 12], but accurate model building requires full state estimation. To address this issue, recent advances in machine learning are applied to learn the system dynamics models directly from sensory data [22, 13, 39]. The learned dynamics model generally demonstrates superior effectiveness relative to its model-free counterparts [19]. Graph-based scene representation processed by GNNs have showcased effectiveness in dynamics modeling due to the spatial relational inductive bias injected [2, 24, 16, 3, 30]. Prior work explored a wide range of applications of graph-based dynamics model, including but not limited to rigid bodies [16, 3], elastic-plastic materials [32, 31], fluids [16, 30], and granular materials [40, 30, 37]. However, the vast majority of prior work model systems that are either confined to a 3D box or atop a 2D plane, and our work demonstrates an approach to extend graph-based approaches to terrains of unbounded breadth and depth using regions of interest and RoI boundary features.
Model-Free Policy Learning. While model-free methods such as reinforcement learning (RL) and behavior cloning (BC) have achieved notable success in robotic manipulation [7, 8, 14, 5, høeg2025hybriddiffusionsimultaneoussymbolic, 18, 6], they typically suffer from limited generalization, particularly in environments involving complex, high-dimensional, and long-horizon interactions like terrain deformation. These approaches require extensive task-specific data collection and tuning, often making them impractical for settings involving diverse or previously unseen terrain geometries. In the context of deformable terrain manipulation, the inability of model-free policies to extrapolate beyond the training distribution presents a fundamental limitation [26]. While such methods can learn reactive strategies within the training distribution, they lack an explicit mechanism for generalizing to new goals or terrain states. In contrast, dynamics model-based approaches retain predictive structure: given a learned forward model, planning can be performed zero-shot for unseen terrain states by simulating the effects of actions. This capability enables flexible and task-agnostic generalization, which is essential for terrain-aware manipulation. Because of these limitations, we do not consider model-free policies as baselines in this work. Our focus is exclusively on dynamics model learning, as it enables both general-purpose planning and scalability across diverse terrain interaction scenarios.
@@ III-C Adaptive Fine-Grained Region of Interest Identification
GBND ’s computational and memory costs scale with the number of particles and edges; therefore, constraining the graph size is critical for efficiency. We learn an implicit function $g(p,u_{t})$ that defines the region of interest (RoI) $R_{t}(u_{t})=\{p\in\mathbb{R}^{3}\mid g(p,u_{t})\geq 0\}$.
@@ III-F Terrain Manipulation Planning with Localized GBND
We use our L-GBND in a planner for terrain manipulation tasks. Our planner selects among parameterized scooping, pushing, and dumping trajectories to minimize the difference between the predicted terrain shape and a target heightmap. For scooping and pushing, we use the penetrate-drag-scoop (PDS) trajectory used in excavation [48, 20, 33], which maintains a 0 roll angle. This yields a 9D motion parameter space. Dumping uses a 6D parameterization (dumping location, yaw angle, pre- and post-dump pitch angles). Scooping and pushing are only allowed if the scoop is empty. The planner is based on Model Predictive Path Integral (MPPI) control [41]. A batch of trajectories is sampled in each iteration. The corresponding control sequences are rolled out with the dynamics model, and the outcomes are scored based on proximity to the goal shape. The best trajectory is selected and executed, and the terrain state is updated. Pseudocode is shown in Algo. 2.